# JingFang: A Traditional Chinese Medicine Large Language Model of Expert-Level Medical Diagnosis and Syndrome Differentiation-Based Treatment

### v1

@@ Appendix C TCM Treatment Recommendation
Appendix D More Information of Ablation Experiment
D.1 Details of the Evaluation of TCM General Agent in Medical Consultation
To evaluate the impact of the combination of TCM Specialist Agent and TCM General Agent in MDCCTM on multi-round medical consultation and to analyze the key role and contributions of the TCM General Agent, this study compares two configurations: JF without TCM General Agent and JF with TCM General Agent, systematically analyzing their differences in performance in multi-round consultation.
@@ D.1 Details of the Evaluation of TCM General Agent in Medical Consultation
D.2 Applicability Evaluation of the TCM Syndrome Differentiation Dataset
@@
Using LoRA, we fine-tuned these models with a training dataset comprising over 43,000 high-quality entries from the TCM syndrome differentiation dataset. The training was conducted on four A6000 GPUs, each with 48GB of VRAM. Key training parameters were set as follows: ‘”lora rank”: 64‘, ‘”lora alpha”: 16‘, ‘”learning rate”: 5e-4‘, and ‘”batch size”: 16‘.
@@
Appendix E Examples of Complete Medical Consultation
@@
E.1 Comprehensiveness of Multi-Round Consultations
@@
E.2 Pertinence of Multi-Round Consultation
@@
E.3 Dynamic Optimization of consultation CoT
@@
E.4 Personalization of Treatment Recommendation
@@
E.5 Conclusion
The application of the JF framework in multi-round TCM consultations fully demonstrates its advantages in comprehensiveness, pertinence, dynamic optimization, and personalized treatment plans. This framework not only enhances the accuracy of TCM diagnosis but also improves the patient experience through the dynamic optimization of the consultation CoT, showcasing significant clinical application value. In the future, this framework can be further integrated with intelligent technologies to enhance the intelligence level of TCM diagnosis and treatment, providing strong support for the development of TCM LLMs.
Figure 12: An example of Damp-Heat Sinking Downward Syndrome
Figure 13: An example of Spleen and Kidney Deficiency Syndrome
Figure 14: An example of Cold Congealing and Blood Stasis Syndrome
Figure 15: An example of Yang Deficiency with Water Overflow Syndrome

### v2

@@ Appendix D TCM Treatment Recommendation
We designed and conducted a comparative experiment to evaluate the performance differences between the DSRS and the direct symptom-based retrieval method. In this study, 100 clinical cases were randomly selected, and both the DSRS and single-stage direct symptom matching were applied under a unified recall threshold. The average similarity to the original data was then compared between the two approaches.
As shown in the Figure 10, the DSRS significantly outperformed the single-stage method in terms of similarity to the original cases. Further analysis revealed that the DSRS effectively reduced semantic noise introduced by non-target syndrome types that share similar symptom keywords, thereby improving the specificity and effectiveness of the retrieval. These findings demonstrate the potential of DSRS in TCM syndrome differentiation and treatment, contributing to more precise and intelligent recommendations in cases of “different treatments for the same disease” or “same treatment for different diseases.”
Figure 10: DSRS Ablation Experiment
Appendix E More Information of Ablation Experiment
E.1 Details of the Evaluation of TCM General Agent in Medical Consultation
To evaluate the impact of the combination of TCM Specialist Agent and TCM General Agent in MACCTM on multi-round medical consultation and to analyze the key role and contributions of the TCM General Agent, this study compares two configurations: JF without TCM General Agent and JF with TCM General Agent, systematically analyzing their differences in performance in multi-round consultation.
@@ E.1 Details of the Evaluation of TCM General Agent in Medical Consultation
E.2 Applicability Evaluation of the TCM Syndrome Differentiation Dataset
@@
Using LoRA, we fine-tuned these models with a training dataset comprising over 43,000 high-quality entries from the TCM syndrome differentiation dataset. The training was conducted on four A6000 GPUs, each with 48GB of VRAM. Key training parameters were set as follows: ‘"lora rank": 64‘, ‘"lora alpha": 16‘, ‘"learning rate": 5e-4‘, and ‘"batch size": 16‘.
@@
Appendix F Examples of Complete Medical Consultation
@@
F.1 Comprehensiveness of Multi-Round Consultations
@@
F.2 Pertinence of Multi-Round Consultation
@@
F.3 Dynamic Optimization of consultation CoT
@@
F.4 Personalization of Treatment Recommendation
@@
F.5 The Performance of JingFang in Authentic Clinical Scenarios
The application of the JF framework in multi-round TCM consultations fully demonstrates its advantages in comprehensiveness, pertinence, dynamic optimization, and personalized treatment plans. This framework not only enhances the accuracy of TCM diagnoses but also improves the syndrome differentiation-based treatment through the systematically designed and developed framework, showing significant value on practical application. Here are some examples of JF’s performance on medical consultation, syndrome differentiation and treatment in authentic clinical scenarios.
Appendix G Limitations
Despite JingFang’s expert-level performance in multi-round TCM consultation and syndrome differentiation, several limitations remain in real-world clinical settings. First, the system primarily relies on textual representations of tongue and pulse diagnostics, lacking the capability to directly process multimodal inputs such as tongue images and pulse waveform signals, thereby limiting its ability to fully emulate the “Four Diagnostic Methods” of TCM. Second, the training data contain substantial noise and irrelevant content, necessitating meticulous data cleaning and alignment to ensure model accuracy. Moreover, real-world consultations are highly complex, dynamic, and patient-specific. Although the current consultation COT framework based on the classical “Ten Questions Song” offers a degree of comprehensiveness and pertinence, it still falls short of fully capturing the diversity of real clinical scenarios. Future work will focus on incorporating multimodal perception mechanisms and developing models with stronger TCM diagnostic and reasoning capabilities to enhance generalizability and practical utility in authentic clinical environments.
Appendix H Broader Impacts
The development of JingFang marks a significant step toward the deep integration of TCM and LLMs, offering promising applications in primary care, remote consultations, and medical education. It also holds potential to facilitate the standardization and intelligent transformation of TCM diagnostic and treatment processes. In long-term health management tasks such as chronic disease management, JingFang can provide continuous and personalized TCM intervention recommendations, supporting patients in achieving more precise and effective recovery. However, the widespread deployment of such models also entails potential risks for instance, overreliance on AI systems may erode physicians’ clinical decision-making capabilities. Therefore, we advocate for a human-centered, assistive-first principle in practical deployment to ensure that the development of TCM-oriented LLMs is grounded in both patient safety and professional medical oversight.
Figure 14: An example of JF’s clinical performance in Damp-Heat Sinking Downward Syndrome differentiation and treatment
Figure 15: An example of JF’s clinical performance in Spleen and Kidney Deficiency Syndrome differentiation and treatment
Figure 16: An example of JF’s clinical performance in Cold Congealing and Blood Stasis Syndrome differentiation and treatment
Figure 17: An example of JF’s clinical performance in Yang Deficiency with Water Overflow Syndrome differentiation and treatment